# Target Speaker ASR with Whisper

### v1

@@ Abstract
We propose a novel approach to enable the use of large, single speaker ASR models, such as Whisper, for target speaker ASR. The key insight of this method is that it is much easier to model relative differences among speakers by learning to condition on frame-level diarization outputs, than to learn the space of all speaker embeddings. We find that adding even a single bias term per diarization output type before the first transformer block can transform single speaker ASR models, into target speaker ASR models. Our target-speaker ASR model can be used for speaker attributed ASR by producing, in sequence, a transcript for each hypothesized speaker in a diarization output. This simplified model for speaker attributed ASR using only a single microphone outperforms cascades of speech separation and diarization by 11% absolute ORC-WER on the NOTSOFAR-1 dataset.
@@ I Introduction
Self-supervised models [1, 2, 3], LLMs [4, 5], and Whisper-style supervised models [6, 7] have demonstrated that scaling up models to use more parameters and extremely large amounts of data can enable the development of accurate automatic speech recognition (ASR) systems, even in relatively challenging environments. However, these models have primarily been used in single-speaker, single-channel ASR systems, whereas most conversations involve multiple talkers and are often recorded by one or more microphones.
@@
Approaches to handle this scenario generally combine multiple systems that perform source separation, speaker segmentation, overlapped speech detection, post-hoc speaker clustering, and ASR in order to produce speaker-attributed conversation transcripts. Alternatively, there are end-to-end systems that transcribe multi-talker speech directly using special tokens or multiple heads [8, 9, 10, 11]. One type of semi-end-to-end system, dubbed target-speaker ASR (TS-ASR) [12, 13, 14, 15, 16], uses the original input mixture and transcribes each speaker separately. Internally, these models rely on source separation to isolate the target speaker’s speech [17, 18].
Another approach to TS-ASR works by learning to transcribe a single target-speaker’s transcript from a mixture of sources by conditioning on a pre-extracted speaker embedding. Typically, this requires a pretrained speaker embedding extractor [19] and training the ASR model from scratch. More recent methods include the use of adaptation layers and soft prompts to modify existing ASR models to work with speaker embeddings [14]. Because these models are often trained on simulated datasets due to the limited availability of multi-talker ASR datasets and the need for a large number of speaker identities, there is signficant performance degradation when such models are deployed on real multi-speaker data “in the wild” [20, 21, 22].
@@
Figure 1: Proposed Diarization-Conditioned Whisper model. An input audio segment with potentially multiple speakers is augmented with frame-level diarization outputs $\begin{bmatrix}p_{S}^{t}&p_{T}^{t}&p_{N}^{t}&p_{O}^{t}\end{bmatrix}^{T}$ for each of the STNO classes at every frame $t$. Affine transformations, indicated as additions to the left of the Whisper model, are applied to intermediate input representations $\mathbf{z}_{n}^{1:T}$ to generate new embeddings, where $n$ stands for the index of the layer. The final frame-level embedding is a convex combination of these embeddings for each frame.
TABLE I: Comparison of the proposed system build with Whisper-large-v3, trained and tested separately on each dataset, alongside various multi-talker ASR systems. The top section includes systems where no additional information about speaker identity or segmentation is provided. The bottom section features models that directly or indirectly utilize segmentation information.
AMI-sdm
NOTSOFAR-1
Libri2Mix
test-both
Kanda et al. [23] 25.8 - -
Raj et al. [8] 44.6 60.9 -
Vinnikov et al. [20] 35.5
Input masking 79.1 76.6 56.7
Proposed 48.5 24.5 17.6
Ma et al. [14] - - 26.4
Zhang et al. [15] - - 23.5
@@
In this paper, we propose a semi-end-to-end approach to TS-ASR that uses Whisper in a new way. Unlike previous TS-ASR methods, our system does not rely on speaker embeddings, but instead conditions directly on frame-level diarization outputs. We believe that, compared to the aforementioned embedding-based approaches, only ”relative” differentiation between speakers is needed; the TS-ASR system does not need to adapt to an existing subspace of speaker embeddings. Training our model on labeled examples of both target and non-target speech, may also improve speaker discrimination and improve robustness to diarization errors.
@@
To validate our approach, we fine-tune Whisper models on the NOTSOFAR-1 [20], AMI [21], and Libri2Mix [24] datasets using ground truth speaker segmentation. During inference, we utilize the ground truth speaker segmentation as well. All experiments conducted in this study adhere to the conditions of the NOTSOFAR-1 Challenge [^1] https://www.chimechallenge.org/current/task2/index.
@@ II-A Silence, Target, Non-Target, and Overlap Masks
Let $\mathbf{D}\in[0,1]^{S\times T}$, where $S$ is the number of speakers in the recording, and $T$ is the number of frames. The matrix $\mathbf{D}$ represents the diarization output, with $d(s,t)$ denoting the probability that speaker $s$ is active at time frame $t$.
@@
The dependency on the number of speakers in $\mathbf{D}$ can be a limiting factor for easily incorporating this mask into the model. To address this, let $s_{k}$ represent the target speaker. We define a distribution over the following mutually exclusive events for a frame at time $t$.
  • S: The time frame, $t$, is silence
  • T: The target speaker, $s_{k}$, is the only active speaker at time frame, $t$.
  • N: One or more non-target speakers, $s\neq s_{k}$ is active and the target speaker, $s_{k}$, is not active at the time frame, $t$.
  • O: The target speaker $s_{k}$ is active while at least one non-target speaker $s\neq s_{k}$ is also active at time frame, $t$.
We define the following distribution over these events.
$p_{S}^{t}=p\left(t=S\right)$ $=\prod_{s=1}^{S}(1-d(s,t))$ (1)
$p_{T}^{t}=p\left(t=T\right)$ $=d(s_{k},t)\cdot\prod_{\begin{subarray}{c}s=1\\ s\neq s_{k}\end{subarray}}^{S}(1-d(s,t))$ (2)
$p_{N}^{t}=p\left(t=N\right)$ $=\left(1-p_{S}^{t}\right)-d\left(s_{k},t\right)$ (3)
$p_{O}^{t}=p\left(t=O\right)$ $=d(s_{k},t)-p_{T}^{t}$ (4)
@@
This definition allows us to use a fix-sized STNO (Silence, Target, Non-target, Overlap) mask $\mathbf{M}^{t}=\begin{bmatrix}p_{S}^{t}&p_{T}^{t}&p_{N}^{t}&p_{O}^{t}\end{bmatrix}^{T}$. The mask is speaker-dependent so that different STNO masks will result in different transcripts by the system.
@@
II-B Input Masking
Having the STNO mask, a straightforward way to perform target speaker ASR is to mask the signal by multiplying each frame by the probability that it is target speech or involves overlap with the target speaker.
Let $\mathbf{X}\in\mathbb{R}^{F\times T}$ denote the matrix of input features, where $F$ is the number of feature dimensions (e.g., mel-filter banks). The masked feature matrix $\mathbf{X}_{\text{masked}}$ is computed as:
$\mathbf{X}_{\text{masked}}(f,t)$ $=\mathbf{X}(f,t)\cdot(p_{T}^{t}+p_{O}^{t}).$ (5)
Here, we add $p_{T}^{t}$ and $p_{O}^{t}$ to ensure that both target speech and overlapping speech are preserved in the masked features. Frames where neither the target speech nor overlap is present are effectively masked out (i.e., set to zero).
However, similar to source separation approaches, this method has limitations. It can introduce artifacts because we are creating a modified version of the input signal, and errors in diarization can propagate through the system, potentially affecting the model’s performance.
II-C Frame-Level Diarization Dependent Transformations
Let $\mathbf{Z}^{n}\in\mathbb{R}^{d_{\text{model}}\times T}$ represent the frame-by-frame inputs to $n$-th (transformer) layer. We transform these states by applying four affine STNO layer- and class-specific transformations $\mathbf{W}_{S}^{n},\mathbf{W}_{T}^{n},\mathbf{W}_{N}^{n},\mathbf{W}_{O}^{n}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ together with biases $\mathbf{b}_{S}^{n},\mathbf{b}_{T}^{n},\mathbf{b}_{N}^{n},\mathbf{b}_{O}^{n}\in\mathbb{R}^{d_{\text{model}}}$ to obtain new speaker-specific states $\hat{\mathbf{Z}}^{n}$ as follows:
$\hat{\mathbf{Z}}^{n}$ $=\left(\mathbf{W}_{S}^{n}\mathbf{Z}^{n}+\mathbf{b}_{S}^{n}\right)\mathbf{p}_{S}+\left(\mathbf{W}_{T}^{n}\mathbf{Z}^{n}+\mathbf{b}_{T}^{n}\right)\mathbf{p}_{T}+$
$+\left(\mathbf{W}_{N}^{n}\mathbf{Z}^{n}+\mathbf{b}_{N}^{n}\right)\mathbf{p}_{N}+\left(\mathbf{W}_{O}^{n}\mathbf{Z}^{n}+\mathbf{b}_{O}^{n}\right)\mathbf{p}_{O}.$ (6)
These transformations create four distinct representations of the frame-by-frame inputs, each emphasizing one of the STNO classes. A new, target-speaker specific representation is formed from the convex combination of these 4 terms, where the term weights come from the STNO mask. Note that the same transformation will be applied to all frames with identical STNO masks.
TABLE III: Effect of enlarging the number of encoder layers, where FDDTs are applied on NOTSOFAR-1 eval-small with Whisper-medium.en (24 encoder layers).
@@ II-C Frame-Level Diarization Dependent Transformations
Including biases in the affine transformations is crucial as it enables the model to differentiate between different STNO types of speech. The biases can shift the representations, making it easier for the model to recognize and distinguish between silence, target speaker speech, non-target speaker speech, and overlapping speech.
On the other hand, by using matrices, $\mathbf{W}_{S,T,N,O}$, we can transform hidden states to the space where it is possible to distinguish between speakers more efficiently or even suppress some parts of the signal.
  • Suppressive Initialization: To bias the model toward masking other speakers, we initialize the $\mathbf{W}_{S,N}$ weights as diagonal matrices with values close to zero, e.g., 0.1. This approach helps the model to distinguish between different types of speech, reinforcing the separation between the STNO classes.
@@
III Experiments
We primarily conducted our experiments on the new NOTSOFAR-1 dataset [20], which includes approximately 315 meetings, each averaging 6 minutes, capturing a broad range of real-world acoustic conditions and conversational dynamics. To verify generalization and demonstrate the competitiveness of the proposed method, we evaluated our best models performance on commonly used synthetic dataset Libri2Mix [24] as well as on real-world meeting recordings like AMI [21] and NOTSOFAR-1. All experiments were conducted in compliance with CHiME8-NOTSOFAR1 rules.
They are divided into two parts. In Section III-C, we examine the behaviour of FDDT under different weight structure constraints, initialization methods, the number of additional parameters, and the information provided. In Section III-D, we analyze the framework’s performance when scaled.
Source codes and recipes [^2] https://github.com/BUTSpeechFIT/TS-ASR-Whisper are built on top of the transformers library [25]. All models are evaluated with the Optimal Reference Combination WER (ORC-WER) [26]. For brevity, we will refer to this metric as WER throughout the rest of the text.
III-A Training details
Adapting the acoustic part of the model without negatively impacting the generalization of the decoder can be challenging, especially with models like Whisper. To address this, we incorporated an additional CTC (Connectionist Temporal Classification) head, following the hybrid CTC-attention-based training scheme proposed in [27]. Given Whisper’s large 50k vocabulary size and fixed sequence length, adding an extra projection layer poses memory challenges. Therefore, we introduced two convolutional layers with a subsampling factor of two each and an additional self-attention layer to optionally realign the sequence. The CTC head can also be used during joint decoding to help mitigate hallucinations [28].
Both the CTC head and the decoder are trained with timestamp tokens. The logits for the CTC blank token are produced by a trainable projection $\mathbf{W}_{\text{blank}}\in\mathbb{R}^{d{{}_{\text{model}}}\times 1}$, and we use $\alpha=0.3$ as the CTC loss weight.
All models are trained with an overall batch size of 64 samples using bf16 precision and the AdamW optimizer [29]. The learning rate is set to $2\times 10^{-6}$, with a weight decay of $1\times 10^{-6}$, a linear decay scheduler, and 2k warm-up steps. The new parameters introduced by FDDT are trained with a learning rate of $2\times 10^{-4}$. By default, FDDT modules are inserted before all layers of the encoder with the diagonal constraint. Unless otherwise stated, the CTC head undergoes an initial ”CTC preheating” phase, where it is trained on LibriSpeech for 10k steps, with the rest of the model frozen. Afterwards, FDDT and CTC parameters are trained for a single epoch (Amplification phase). Finally, the full model is trained for up to 50k steps, with early stopping set to patience of 5 epochs. Most of the models typically converge within ten epochs. For the final evaluation, we always select the best-performing checkpoint based on the development set WER.
TABLE V: Different sizes of training corpuses affecting the performance of Whisper-medium.en. Tested on NOTSOFAR-1 Eval Small.
@@ III-A Training details
III-B Baseline Comparison
Table I compares the proposed method with different end-to-end and modular systems. It can be seen that our approach vastly outperforms the naive approach input masking on all three datasets mainly due to finetuning and the ability to handle overlapped speech. Even though our approach does not outperform any baselines in the first part of the table, it out performs the NOTSOFAR baseline [20] and the finetuned SURT model on the NOTSOFAR-1 dataset as well. Lastly, our approach outperforms all other approaches on the synthetic dataset Libri2Mix.
III-C Frame-Level Diarization Dependent Transformation
To evaluate the impact of the Frame-by-Frame Diagonal Transformations (FDDT), we conducted a study to determine whether limiting the number of parameters in the additional modules affects system performance and how important correct initialization is. Table II shows that using biases alone performs similarly to diagonal matrices, indicating that biasing frame-by-frame representations can effectively focus the model on frames corresponding to the same class. The table also highlights that randomly initializing FDDT parameters is suboptimal and can significantly harm model performance, suggesting that suppressive initialization is preferable. Interestingly, using full weights leads to noticeable performance degradation, likely because it disrupts the frame-by-frame representations. Table III demonstrates that even a single layer of bias-only parameters can achieve performance comparable to the best diagonal setup.
Table IV further shows that a model can perform TS-ASR with just a single STNO class corresponding to the target frames without significant performance degradation. Interestingly, when using three classes (TNO), the model performs worse than when using STNO or TN, which is counterintuitive since the model theoretically receives the same amount of information.

### v2

@@ Abstract
We propose a novel approach to enable the use of large, single-speaker ASR models, such as Whisper, for target speaker ASR. The key claim of this method is that it is much easier to model relative differences among speakers by learning to condition on frame-level diarization outputs than to learn the space of all speaker embeddings. We find that adding even a single bias term per diarization output type before the first transformer block can transform single-speaker ASR models into target-speaker ASR models. Our approach also supports speaker-attributed ASR by sequentially generating transcripts for each speaker in a diarization output. This simplified method outperforms baseline speech separation and diarization cascade by 12.9 % absolute ORC-WER on the NOTSOFAR-1 dataset.
@@ I Introduction
Self-supervised models [1, 2, 3], LLMs [4, 5], and Whisper-style supervised models [6, 7] have demonstrated that scaling up models by using more parameters and extremely large amounts of data can enable the development of accurate automatic speech recognition (ASR) systems, even in relatively challenging environments. However, these models have primarily been used in single-speaker, single-channel ASR systems, whereas most conversations involve multiple speakers and are often recorded by one or more microphones.
@@
Approaches to handling this scenario generally integrate multiple blocks performing source separation, speaker segmentation, overlapped speech detection, post-hoc speaker clustering, and ASR in order to produce speaker-attributed conversation transcripts. Alternatively, there are end-to-end systems transcribing multi-speaker speech directly using special tokens or multiple heads [8, 9, 10, 11]. Another approach is the use of a semi-end-to-end system, known as target-speaker ASR (TS-ASR) [12, 13, 14, 15, 16, 17] processing the original input mixture and transcribing each speaker individually.
The conventional TS-ASR framework extracts speaker embeddings corresponding to the target speakers and uses them as an auxiliary input to the ASR system [18, 16]. While pre-trained speaker embedding extractors [19, 20, 21] can guide the ASR system by highlighting relevant information in the input and filtering out irrelevant content, they inherently require the system to learn how to map speaker embeddings to ASR speech embeddings. More recent methods include the use of adaptation layers and soft prompts to modify existing ASR models to incorporate speaker embeddings [14] or speaker enrollment audios [17]. However, since these models are often trained on simulated datasets—due to the scarcity of multi-speaker ASR datasets and the need for a large number of speaker identities—they tend to experience significant performance degradation when applied to real-world multi-speaker scenarios [22, 23, 24].
@@
Figure 1: Proposed Diarization-Conditioned Whisper model: An input audio segment with multiple speakers is conditioned by frame-level diarization outputs $\begin{bmatrix}p_{\mathcal{S}}^{t}&p_{\mathcal{T}}^{t}&p_{\mathcal{N}}^{t}&p_{\mathcal{O}}^{t}\end{bmatrix}^{T}$ for each STNO class at every frame $t$. Affine transformations, applied to intermediate input representations $\mathbf{z}^{l}_{1:T}$, generate new embeddings, where $l$ is the layer index. The final frame-level embedding is a convex combination of these embeddings for each frame.
@@
In this paper, we propose a semi-end-to-end approach to TS-ASR that uses Whisper in a new way. Unlike previous TS-ASR methods, our system does not rely on speaker embeddings, but instead conditions directly on frame-level diarization outputs. We believe that, compared to the aforementioned embedding-based approaches, only ”relative” differentiation between speakers is needed; the TS-ASR system does not need to adapt to an existing subspace of speaker embeddings. Training our model on labeled examples of both target and non-target speech may also improve speaker discrimination and improve robustness to diarization errors.
@@
To validate our approach, we fine-tune Whisper models on the NOTSOFAR-1 [22], AMI [23], and Libri2Mix [25] datasets using ground truth speaker segmentation. Unless stated otherwise, ground truth segmentation is also used during inference. All experiments follow the NOTSOFAR-1 Challenge guidelines [^1] https://www.chimechallenge.org/challenges/chime8/task2/.
@@ II-A Silence, Target, Non-Target, and Overlap Masks
Let $\mathbf{D}\in[0,1]^{S\times T}$, where $S$ is the number of speakers in the recording, and $T$ is the number of frames, represent the diarization output, with $d(s,t)$ denoting the probability that speaker $s$ is active at time frame $t$.
@@
The dependency on the number of speakers in $\mathbf{D}$ can be a limiting factor for easily incorporating this mask into the model. To address this, let $s_{k}$ represent the target speaker. We define a distribution over the following mutually exclusive events for a frame at time $t$:
  • ${\mathcal{S}}$: Silence at time $t$.
  • ${\mathcal{T}}$: Only the target speaker $s_{k}$ is active at $t$.
  • ${\mathcal{N}}$: Non-target speaker(s) ($s\neq s_{k}$) active; target $s_{k}$ inactive at $t$.
  • ${\mathcal{O}}$: Target $s_{k}$ and at least one non-target ($s\neq s_{k}$) active, causing overlap at $t$.
The probabilities of these events occurring at time frame $t$ can be calculated as:
$p_{\mathcal{S}}^{t}$ $=\prod_{s=1}^{S}(1-d(s,t)),\quad p_{\mathcal{T}}^{t}=d(s_{k},t)\prod_{\begin{subarray}{c}s=1\\ s\neq s_{k}\end{subarray}}^{S}(1-d(s,t))$
$p_{\mathcal{N}}^{t}$ $=\left(1-p_{\mathcal{S}}^{t}\right)-d\left(s_{k},t\right),\quad p_{\mathcal{O}}^{t}=d(s_{k},t)-p_{\mathcal{T}}^{t}$ (1)
@@
This definition allows us to use a fixed-sized target-speaker-dependent STNO (Silence, Target, Non-target, Overlap) mask $\begin{bmatrix}p_{\mathcal{S}}^{t}&p_{\mathcal{T}}^{t}&p_{\mathcal{N}}^{t}&p_{\mathcal{O}}^{t}\end{bmatrix}^{\top}$.
II-B Input Masking
Having the STNO mask, a straightforward way to perform target speaker ASR is to mask the signal by multiplying each frame by the probability that it is target speech or that it involves overlap with the target speaker. Hence, if the target speaker is not active, the audio signal is set to 0 (i.e., silence). We add $p_{\mathcal{T}}^{t}$ and $p_{\mathcal{O}}^{t}$ to ensure that both target speech and overlapping speech are preserved in the masked signal. However, similar to source separation approaches, this method has limitations. It can introduce artifacts because we are creating a modified version of the input signal, and errors in diarization can propagate through the system, potentially affecting the model’s performance.
II-C Frame-Level Diarization Dependent Transformations
Let $\mathbf{Z}^{l}\in\mathbb{R}^{d_{{m}}\times T}$ represent the frame-by-frame inputs to the $l$-th (Transformer) layer. We transform these hidden representations by applying four affine STNO layer- and class-specific transformations: $\mathbf{W}_{\mathcal{S}}^{l},\mathbf{W}_{\mathcal{T}}^{l},\mathbf{W}_{\mathcal{N}}^{l},\mathbf{W}_{\mathcal{O}}^{l}\in\mathbb{R}^{d_{{m}}\times d_{{m}}}$ together with biases $\mathbf{b}_{\mathcal{S}}^{l},\mathbf{b}_{\mathcal{T}}^{l},\mathbf{b}_{\mathcal{N}}^{l},\mathbf{b}_{\mathcal{O}}^{l}\in\mathbb{R}^{d_{m}}$ to obtain new speaker-specific hidden representations $\hat{\mathbf{Z}}^{l}=[\hat{\mathbf{z}}^{l}_{1},\ldots,\hat{\mathbf{z}}^{l}_{T}]$ as
$\hat{\mathbf{z}}^{l}_{t}=$ $\left(\mathbf{W}_{\mathcal{S}}^{l}\mathbf{z}^{l}_{t}+\mathbf{b}_{\mathcal{S}}^{l}\right)p^{t}_{\mathcal{S}}+\left(\mathbf{W}_{\mathcal{T}}^{l}\mathbf{z}^{l}_{t}+\mathbf{b}_{\mathcal{T}}^{l}\right)p^{t}_{\mathcal{T}}$
$+\left(\mathbf{W}_{\mathcal{N}}^{l}\mathbf{z}^{l}_{t}+\mathbf{b}_{\mathcal{N}}^{l}\right)p^{t}_{\mathcal{N}}+\left(\mathbf{W}_{\mathcal{O}}^{l}\mathbf{z}^{l}_{t}+\mathbf{b}_{\mathcal{O}}^{l}\right)p^{t}_{\mathcal{O}},$ (2)
These transformations generate four distinct representations of the frame-by-frame inputs, each highlighting one of the STNO classes. A target-speaker-specific representation is formed by a convex combination of these terms, with weights derived from the STNO mask. The same transformation is applied to all frames with identical STNO masks.
  • Suppressive Initialization: To bias the model toward masking other speakers, we initialize the $\mathbf{W}_{\mathcal{S}}^{l},\mathbf{W}_{\mathcal{N}}^{l}$ weights as diagonal matrices with values close to zero, e.g., 0.1. This approach helps the model to distinguish between different types of speech, reinforcing the separation between the STNO classes.
III Experiments
We primarily conducted our experiments on the new NOTSOFAR-1 dataset [22], which includes 280 meetings, each averaging 6 minutes, capturing diverse real-world acoustic conditions and conversational dynamics. To assess generalization and competitiveness, we also evaluated our best models on the synthetic Libri2Mix dataset [25] and on real-world meeting dataset AMI [23]. Experiments are divided into two parts: In Section III-C, we analyze the FDDT behavior under different weight structure constraints, initializations, numbers of additional parameters, and provided information. In Section III-D, we evaluate the framework’s performance as more parameters and data are added.
Source codes and recipes [^2] https://github.com/BUTSpeechFIT/TS-ASR-Whisper are built on top of the transformers library [26]. All models are evaluated with the Time-Constrained Optimal Reference Combination Word Error Rate (tcORC WER) [27], referred to as WER throughout the text.
TABLE I: Comparison of the proposed system built with updated Whisper-large-v3-turbo [28] alongside various multi-talker ASR systems. The top section includes systems where no additional ground truth information about speaker identity or segmentation is provided. The bottom section features models that directly or indirectly utilize ground truth segmentation information. Results marked with $\dagger$ are evaluated on utterance groups. Proposed ORC WER results marked with $\star$ were approximated by increasing the collar for tcORC WER.
AMI-sdm
test
NOTSOFAR-1
eval-small
Libri2Mix
test-both
ORC WER
Raj et al. [8] $44.6^{\dagger}$ $60.9^{\dagger}$
Vinnikov et al. [22] $35.5$
Niu et al. [29] 17.7
Proposed $18.0^{\star}$ $22.6$ 14.9
Input masking 79.1 76.6 56.7
Ma et al. [14] 26.4
Zhang et al. [15] 23.5
Proposed $16.5^{\star}$ $19.1$ 10.9
III-A Training details
To enhance Whisper’s performance, we incorporated an additional CTC (Connectionist Temporal Classification) head, following the hybrid CTC-attention-based training scheme proposed in [30]. Given Whisper’s large 50k vocabulary size and fixed sequence length, adding an extra projection layer poses memory challenges. To overcome this, we added two convolutional layers, each with a subsampling factor of two, along with an additional self-attention layer. Both the CTC head and the decoder are trained with timestamp tokens, and the CTC loss weight is set to 0.3.
For the ablation experiments, we used Whisper-medium.en, while the final model was trained with Whisper-large-v3-turbo. All models are trained with an overall batch size of 64 samples using bf16 precision and the AdamW optimizer [31]. The learning rate is set to $2\times 10^{-6}$, with a weight decay of $1\times 10^{-6}$, a linear decay scheduler, and 2k warm-up steps. The new parameters introduced by FDDT are trained with a learning rate of $2\times 10^{-4}$. By default, FDDT modules are inserted before all layers of the encoder with the diagonal constraint, meaning that only the diagonal values of the weight matrices are updated during training and can be non-zero. Unless otherwise stated, the CTC head undergoes an initial “CTC preheating” phase, where it is trained on LibriSpeech for 10k steps, with the rest of the model being frozen. Afterwards, FDDT and CTC parameters are trained for a single epoch (FDDT preheating). Finally, the full model is trained for up to 50k steps, with early stopping set to a patience of 5 epochs. Most of the models typically converge within ten epochs. For the final evaluation, we always select the best-performing checkpoint based on the development set WER.
III-B Comparison to Baselines
Table III presents a comparison of the proposed method with various end-to-end and modular systems. The top section includes systems where no additional ground truth information about speaker identity or segmentation is provided, and the system must infer this information. For the proposed system, we utilized DiariZen [32] [^3] https://github.com/BUTSpeechFIT/DiariZen to condition our model. The bottom section features models that directly or indirectly utilize ground truth segmentation information. It can be seen that our approach significantly outperforms the naive input masking baseline across all three datasets, largely due to its fine-tuning capabilities and robust handling of overlapped speech. Although our method does not achieve the top performance on the NOTSOFAR dataset, it surpasses baseline models, including the NOTSOFAR baseline [22] and the fine-tuned SURT model [8]. Furthermore, our approach delivers strong results on the AMI and Libri2Mix datasets, underscoring its effectiveness across diverse scenarios.
III-C Frame-Level Diarization Dependent Transformation
@@
To assess the impact of FDDT, we evaluated whether reducing the number of parameters in the additional modules affects the performance and examined the importance of proper initialization. As shown in Table II, using biases alone performs comparably to the combination of diagonal transformation matrices and biases, indicating that biasing frame-by-frame representations is sufficient to effectively focus the model on frames belonging to the same STNO class. The table also highlights that randomly initializing FDDT parameters is suboptimal and can significantly harm model performance, suggesting that suppressive initialization is preferable. Interestingly, using non-restricted weights, meaning full matrices without diagonal constraints leads to noticeable performance degradation.
TABLE III: Effect of increasing the number of encoder layers, where FDDTs are applied on NOTSOFAR-1 eval-small with Whisper-medium.en (24 encoder layers).
@@
Table III demonstrates the effect of increasing the number of encoder layers where FDDTs are applied. The number of layers refers to how many layers, starting from the first encoder layer, are modified by FDDT parameters. For example, when only one layer is used, only the input to the Whisper encoder is modified. The results indicate that even a single layer of bias-only parameters can achieve performance comparable to the best diagonal setup. However, using a random initialization with diagonal matrices, even for just the first layer, severely impacts the model’s performance. Notably, the use of a diagonal setup with both $\mathbf{W}_{diag}$ and $\mathbf{b}$ across 24 layers yields the best performance, with the suppressive initialization method achieving the lowest WER of 26.7.
@@
Table IV presents the performance of the FDDT method under different reductions of diarization information. It demonstrates that while the FDDT model still significantly outperforms input masking, some performance degradation is evident when comparing the STNO and T configurations. Interestingly, using three classes (TNO) results in worse performance than using (TN) classes only, even though the STNO and TNO configurations provide the same amount of information.
TABLE V: Different sizes of training corpora affecting the performance of Whisper-medium.en. Tested on NOTSOFAR-1 Eval Small.